OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (6, 2) and has a slope of 3? y = x +

OpenStudy (anonymous):

We can first plug it into point-slope form: \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value of the point, \(x_1\) is the x-value of the point, and \(m\) is the slope. So we have: \(y - 2 = 3(x - 6)\) Now distribute 3 into the parenthesis: \(y - 2 = 3x - 18\) Now add 2 to both sides, what's -18 + 2?

OpenStudy (anonymous):

y=3x+-4

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Add -18 + 2

OpenStudy (anonymous):

-16

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Yes, so we have: \(y = 3x - 16\) So '3' goes in the first blank and '-16' goes in the 2nd.

OpenStudy (anonymous):

\(\bbox [10pt, lime, border:5pt solid black]{\Huge\cal\color{red}\diamondsuit\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\diamondsuit}}}\) You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

Thanks can you help with more

OpenStudy (anonymous):

Sure, just close this one and open a new one.